OpenStudy (anonymous):

The table below represents the distance of a car from its destination as a function of time. Part A: What is the y-intercept of the function and what does this tell you about the car? (4 points) Part B: Calculate the average rate of change of the function represented by the table between x = 1 to x = 3 and what does the average rate represent? (4 points) Part C: What would be the domain of this function if the car does not change motion?

OpenStudy (anonymous):

Time Distance (Hour) (miles) x y 0 50 1 52 2 54 3 56

OpenStudy (timmyk14):

The y-intercept is 50

OpenStudy (timmyk14):

That tells us that the initial distance the car traveled at hour 0 is 50 miles.

OpenStudy (timmyk14):

Average rate of change is done using this formula:\[\frac{ f(b) - f(a) }{ b - a }\] Where f(b) and f(a) represent the output values, or in this case the miles the car has traveled. The b and the a represent the hour, or the input values.

OpenStudy (timmyk14):

So this is how it would look for this scenario: \[\frac{ f(3) - f(1) }{ 3 - 1 }\] This also means: \[\frac{ 56 - 52 }{ 3 - 1 }\]

OpenStudy (timmyk14):

Before we proceed I would like some feedback from you. What is the average rate of change between hour 1 to hour 3? All you have to do is solve the equation I have you above.

OpenStudy (anonymous):

It is 2

OpenStudy (timmyk14):

Right, the average rate of change from hour 1 to hour 3 is 2. This means that, on average, the car traveled 2 miles every hour from hour 1 to hour 3.

OpenStudy (timmyk14):

I believe that the domain for this function would be \[0\le x \le3\] **NOTE: I said I believe, which means that I'm not sure. This is the same thing I'm currently learning. I'm 90% positive that I'm right though.**

OpenStudy (timmyk14):

Medals would be appreciated...and you're welcome for the help

OpenStudy (anonymous):

You are awesome thank you.